5. The 50 American states have areas (in thousands of square miles) that range from 591 (Alaska) to 1.2 (Rhode Island). The median is 56.2 and the mean is 72.4. If possible, compute the total area. If not, say which further information you would need.
Answer #1

Ans

total area = Mean *number of states = 72.4*50 = 3620 (in thousands of square miles)